OpenStudy (anonymous):

If c is equal to the sum of a and twice b, which of the following is the average (arithmetic mean) of a and c? (a) c (b) a + b (c) a (d) b (e) a +c

OpenStudy (helder_edwin):

c=a+2b

OpenStudy (helder_edwin):

The average of a and c would be \[ \large \frac{a+c}{2}= \]

OpenStudy (anonymous):

yeah how do i solve that

OpenStudy (helder_edwin):

just plug in the value of c

OpenStudy (helder_edwin):

\[ \large \frac{a+c}{2}=\frac{a+(a+2b)}{2}= \]

OpenStudy (anonymous):

oh is see is it (b) a + b?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

ok thank you